OpenStudy (anonymous):

what is the simplified form of x^2-16/x+4 ?

OpenStudy (anonymous):

Zero denominator expression makes the expression undefined [even though it's reducible].. x + 4 = 0 x = -4 If we factor x² - 16, then we obtain: (x - 4)(x + 4)/(x + 4) = x - 4 Hence, the answer is x – 4,

OpenStudy (anonymous):

whats the restriction of this problem?

OpenStudy (anonymous):

i dont know :l

jimthompson5910 (jim_thompson5910):

The restriction is that x can't equal -4

jimthompson5910 (jim_thompson5910):

This is so the original expression and the result are equivalent

OpenStudy (anonymous):

thankkkkksssss. :)